# An Improved Algorithm for Adversarial Linear Contextual Bandits via Reduction

### v1

@@ 3.4 Robust Linear Contextual Bandits
As discussed in Section 3.3, we would like to develop a linear bandit algorithm that tolerates misspecification. Although there is a rich related literature, most prior results are for the stochastic linear bandit problem and do not apply here. For adversarial linear bandits that is misspecification robust, we are only aware of the work by Neu and Olkhovskaya, (2020) and Liu et al., 2024a . However, the bound in Neu and Olkhovskaya, (2020) has a worse $T^{2/3}$ regret, while the algorithms of Liu et al., 2024a are either computationally inefficient or highly sub-optimal. Fortunately, our problem is slightly easier than that studied by Liu et al., 2024a , as our learner has knowledge of the amount of misspecification $\epsilon$, and this amount remains the same in all rounds. This allows us to design the computationally efficient Algorithm 2 with an improved dependence on the amount of misspecification.
Algorithm 2 is an adaptation of the clipped continuous exponential weight algorithm of Ito et al., (2020): The $q_{t}$ in Line 2 of Algorithm 2 is the standard continuous exponential weights, while the $\widehat{q}_{t}$ in Line 2 confines the support of $q_{t}$ within an ellipsoid centered around the mean. This is helpful in obtaining a first-order bound (Ito et al.,, 2020). Sampling from $\widehat{q}_{t}$ can be done with standard techniques for sampling from a log-concave distribution (to sample from $q_{t}$) plus rejection sampling (to correct the distribution to $\widehat{q}_{t}$), which admits a $\mathop{\mbox{poly}}(d,C,T)$ computational complexity. See Ito et al., (2020)’s Section 4.4 for a discussion on the computational complexity.
The key addition compared to (Ito et al.,, 2020) is the bonus term $b_{t}$ that ensures misspecification robustness. This bonus encourages additional exploration, preventing the learner from being misled by misspecified feedback. The form of the bonus for adversarial linear bandits was first developed in the series of work (Lee et al.,, 2020; Zimmert and Lattimore,, 2022) aimed at obtaining high-probability bounds. Our use of the bonus is similar to Liu et al., 2024b , which tackles corruption and misspecification. In the regret analysis, this bonus term creates a negative regret term that offsets the regret overhead due to misspecification.
@@
We remark that while there exist algorithms that are $1$-misspecification-robust (Liu et al., 2024a, ), their run time scales at least with the number of actions. Algorithm 2 achieves $\alpha$-robustness with the smallest $\alpha$ we are aware of among algorithms that run in $\mathop{\mbox{poly}}(d,C,T)$ time.
@@
1 Parameters: $\gamma=10\log(10dT)$, $\beta=T^{-4}$.
@@ 3   Define
$q_{t}(y)=\frac{\exp\left(-\eta\sum_{\tau<t}\left\langle y,\widehat{\theta}_{\tau}-b_{\tau}\right\rangle\right)}{\int_{\widehat{\Omega}}\exp\left(-\eta\sum_{\tau<t}\left\langle z,\widehat{\theta}_{\tau}-b_{\tau}\right\rangle\right)\mathrm{d}z},\quad x_{t}=\mathbb{E}_{y\sim q_{t}}[y],\quad\Sigma_{t}=\mathbb{E}_{y\sim q_{t}}[(y-x_{t})(y-x_{t})^{\top}].$
@@ 4   Define
$\widehat{q}_{t}(y)=\frac{q_{t}(y)\mathbb{I}\left\{\|y-x_{t}\|_{\Sigma_{t}^{-1}}^{2}\leq d\gamma^{2}\right\}}{\int_{\widehat{\Omega}}q_{t}(z)\mathbb{I}\left\{\|z-x_{t}\|_{\Sigma_{t}^{-1}}^{2}\leq d\gamma^{2}\right\}\mathrm{d}z},\quad\widehat{\Sigma}_{t}=\mathbb{E}_{y\sim\widehat{q}_{t}}[(y-x_{t})(y-x_{t})^{\top}].$
@@
6  [3pt] Define $\widehat{\theta}_{t}=(\beta I+\widehat{\Sigma}_{t})^{-1}(y_{t}-x_{t})c_{t}$ and
$b_{t}=8\eta\left(\epsilon+\frac{1}{T^{2}}\right)\sum_{\tau<t}(\widehat{\theta}_{\tau}-b_{\tau}).$
@@ 3.5 Combining Everything and Using the Doubling Trick
Given access to a $\alpha$-misspecification-robust adversarial linear bandit algorithm $\mathsf{ALG}$, Algorithm 1 with doubling trick achieves $\min_{\pi\in\Pi}\mathrm{Reg}_{T}(\pi)\leq\widetilde{O}(d\sqrt{T}+\alpha d\sqrt{T\log K})$ in adversarial linear contextual bandits without access to simulators.
@@
$\min_{\pi\in\Pi}\mathrm{Reg}_{T}(\pi)\leq\tilde{O}\left(\sum_{k=1}^{\log_{2}T}d\sqrt{2^{k}}+\sum_{k=1}^{\log_{2}T}\alpha d\sqrt{2^{k}\log(NKT)}\right)=\tilde{O}\left(d\sqrt{T}+\alpha d\sqrt{T\log K}\right).$
@@ 4 Computational Complexity
Algorithm 1 contains two steps, lines 1 and 1, for which it is not obvious whether they can be implemented efficiently. We will now describe an approach to implement both steps in $\mathop{\mbox{poly}}(d,C,N)$ time, provided that $\operatorname{conv}(\mathcal{A})$ is a polytope that can be described by at most $C$ linear constraints for $\mathcal{D}$-almost all $\mathcal{A}$.
We start with the computational complexity of a separation oracle for $\widehat{\Omega}$. We assume, without loss of generality, that no two sets $\mathcal{A}_{i}$ and $\mathcal{A}_{j}$ in the construction of $\widehat{\Omega}$ are equal; otherwise we can replace them by a single set $2\mathcal{A}_{i}$, which will only decrease $N$. Thus, $\widehat{\Omega}$ is the Minkowski sum of $N$ convex polytopes that, by assumption, can all be described by at most $C$ constraints. For any $x\in\mathbb{R}^{d}$, let
@@
If $g(\phi)>0$ for some $\phi$, then $\phi$ gives a hyperplane that separates $x$ from $\widehat{\Omega}$; and if $g(\phi)\leq 0$ for all $\phi$, then $x\in\widehat{\Omega}$. Since $g$ is concave, and we can solve every subproblem $\max_{x_{i}\in\mathcal{A}_{i}}\phi^{\top}x_{i}$ in $\mathop{\mbox{poly}}(d,C)$ time, we can maximize $g$ in $\mathop{\mbox{poly}}(d,N,C)$ time to obtain our separation oracle.
@@
We then proceed with line 1 of Algorithm 1. To implement it, we need to be able to take any point $y_{t}\in\widehat{\Omega}$ and find a distribution $\alpha_{t}$ supported on the vertices of $\widehat{\Omega}$ that we can sample from efficiently. By Carathéodory’s theorem, any $y_{t}\in\widehat{\Omega}$ can be represented as a convex combination of at most $d+1$ vertices of $\widehat{\Omega}$: $y_{t}=\sum_{l=1}^{k}\alpha_{l}v_{l}$ for $k\leq d+1$ where all $v_{l}$ are vertices of $\widehat{\Omega}$, and $\alpha_{1},\dots,\alpha_{k}\geq 0$ with $\sum_{l}\alpha_{l}=1$ can be interpreted as the probabilities of selecting the vertices. This is a categorical distribution on at most $d+1$ points, from which we can sample in $O(d)$ time (assuming we have access to an oracle that provides samples from the uniform distribution on $[0,1]$). Thus the main challenge is to compute the vertices $v_{l}$ and the probabilities $\alpha_{l}$. By Corollary 14.1g) of Schrijver, (1986) (restated in Lemma 3), there exists an algorithm that can do both in time $\mathop{\mbox{poly}}(d,h)$ for any set $\widehat{\Omega}$ given access to a separation oracle that runs in time $O(h)$. As discussed above, we have $h$ of order $\mathop{\mbox{poly}}(d,C,N)$, from which it follows that we can also find $\alpha_{l}$ and sample from it in $\mathop{\mbox{poly}}(d,C,N)$ time.
[…]
By Corollary 14.1a) of Schrijver, (1986) (restated in Lemma 4) and assuming access to a separation oracle for $\widehat{\Omega}$, we can compute $x^{\star}$ in polynomial time. This gives a separation oracle for the feasible region defined by the strict cone constraint. As a result, we can use the ellipsoid method (e.g., Grötschel et al., (1988), Chapter 3) to find a feasible $\phi$ in polynomial time. The total runtime is $\mathop{\mbox{poly}}(d,C,N)$.
@@ 5 Small-Loss Bound with Access to Simulator
A misspecification-robust linear bandit algorithm with small-loss bounds over action set $\widehat{\Omega}\subset\mathbb{R}^{d}$ has the following property: with a given $\epsilon>0$ and the guarantee that every time the learner chooses $y_{t}\in\widehat{\Omega}$, the loss received $c_{t}\in[0,1]$ satisfies
the algorithm ensures
@@ 6 Conclusion and Open Questions
Three open questions are left by our work: First, can the computation cost be improved further to match that of Neu and Valko, (2014), who only require a linear optimization oracle for each action set, and do not require a polynomial number of constraints? Second, can the regret be further improved to the near-optimal $\widetilde{\mathcal{O}}(d\sqrt{T})$ bound with a polynomial-time algorithm without simulators? Third, can we achieve $\mathop{\mbox{poly}}(d)\sqrt{L^{\star}}$ regret without simulator? For the second question, one may try to generalize the approach of Schneider and Zimmert, (2023). For the third question, one idea is to establish a Bernstein-type counterpart of Lemma 1, potentially drawing ideas from Bartlett et al., (2005); Liang et al., (2015).
We expect that our approach has wider applications than adversarial linear contextual bandits. For example, our approach may be generalized to linear MDPs with fixed transition and adversarial losses (Luo et al.,, 2021; Sherman et al.,, 2023; Dai et al.,, 2023; Kong et al.,, 2023; Liu et al., 2024b, ) and facilitate learning with exponentially large or continuous action sets.
Appendix A Comparison with Hanna et al., (2023)
Our work and Hanna et al., (2023) both reduce linear contextual bandits to linear bandits with misspecification. The linear bandits the two works reduce to are slightly different. For comparison, let $\mathcal{A}_{1},\ldots,\mathcal{A}_{N}$ be empirical action sets drawn from $\mathcal{D}$, and define
@@
It holds that $\widehat{\Omega}_{\text{lin}}^{\epsilon}\subset\widehat{\Omega}_{\text{lin}}\subset\widehat{\Omega}$ because $\Pi^{\epsilon}_{\text{lin}}\subset\Pi_{\text{lin}}\subset\Pi$. The relation among the three sets in (6) is the following: $\widehat{\Omega}_{\text{lin}}$ is the boundary of $\widehat{\Omega}$ (Lemma 2), and $\widehat{\Omega}_{\text{lin}}^{\epsilon}$ is a discretized subset of $\widehat{\Omega}_{\text{lin}}$. It has been shown that there must exist an optimal policy in $\Pi_{\text{lin}}$ (see our arguments in Section 2.1), and that $\Pi_{\text{lin}}^{\epsilon}$ must contain an $\epsilon$-near-optimal policy (see (47) of Liu et al., (2023)).
Our work reduces the linear contextual bandit problem to linear bandits in $\widehat{\Omega}$, while Hanna et al., (2023) reduces it to linear bandits in $\widehat{\Omega}^{\epsilon}_{\text{lin}}$. According to the discussion above, both suffice to establish no regret guarantees against the optimal policy. The key difference lies in the computational efficiency. As discussed in Section 4, $\widehat{\Omega}$ is a polytope whose separation oracle can be implemented in $\mathop{\mbox{poly}}(d,C,T)$ time, while in Hanna et al., (2023), $\widehat{\Omega}^{\epsilon}_{\text{lin}}$ is a set containing $(\frac{1}{\epsilon})^{\Theta(d)}$ discrete points with $\epsilon$ chosen as $\Theta(\frac{1}{T})$. It remains open whether $\widehat{\Omega}^{\epsilon}_{\text{lin}}$ can be represented efficiently and whether Hanna et al., (2023)’s algorithm can be implemented in polynomial time. In addition, our work operates on the adversarial-loss setting, strictly generalizing Hanna et al., (2023)’s stochastic-loss setting.
A technical question related to the discussion above is whether we can apply the discretization idea from Hanna et al., (2023) purely in the analysis but not in the algorithm, i.e., bounding the regret overhead by the discretization error without restricting the learner’s policy to the discretized linear policies. This would allow us to avoid the Rademacher complexity analysis (Lemma 1) and instead use Hoeffding’s or Bernstein’s inequality combined with a union bound, possibly leading to an improved regret bound. However, there is a key technical challenge: the mapping $\hat{\Psi}(\pi_{\phi})$ is not continuous in $\phi$, which prevents us from bounding $\|\hat{\Psi}(\pi_{\phi})-\hat{\Psi}(\pi_{\phi^{\prime}})\|$ by a constant times $\|\phi-\phi^{\prime}\|$, where $\phi$ is an arbitrary point and $\phi^{\prime}$ is a point in the discretized set. This would leave us no control on the performance of policies outside the set of discretized linear policies $\Pi^{\epsilon}_{\text{lin}}$. As a result, we are unable to derive a uniform convergence result analogous to Lemma 1 using this approach.

### v2

@@ 3.4 Robust Linear Contextual Bandits
As discussed in Section 3.3, we would like to develop a linear bandit algorithm that tolerates misspecification. Although there is a rich related literature, most prior results are for the stochastic linear bandit problem and do not apply here. For adversarial linear bandits with misspecification robustness, we are only aware of the work by neu20 and liucorruption. However, the bound in neu20 has a worse $T^{2/3}$ regret, while the algorithms of liucorruption are either computationally inefficient or highly sub-optimal. Fortunately, our problem is slightly easier than that studied by liucorruption, as our learner has knowledge of the amount of misspecification $\epsilon$, and this amount remains the same in all rounds. This allows us to design the computationally efficient Algorithm 2 with an improved dependence on the amount of misspecification.
Algorithm 2 is an adaptation of the clipped continuous exponential weight algorithm of ito2020tight: The $q_{t}$ in Line 2 of Algorithm 2 is the standard continuous exponential weights, while the $\widehat{q}_{t}$ in Line 2 confines the support of $q_{t}$ within an ellipsoid centered around the mean. This is helpful in obtaining a first-order bound (ito2020tight). Sampling from $\widehat{q}_{t}$ can be done with standard techniques for sampling from a log-concave distribution (to sample from $q_{t}$) plus rejection sampling (to correct the distribution to $\widehat{q}_{t}$), which admits a $\mathop{\mbox{poly}}(d,C,T)$ computational complexity, provided that we have access to an efficient linear optimization oracle for $\hat{\Omega}$. See ito2020tight’s Section 4.4 for a discussion on the computational complexity. We discuss how to construct such an efficient linear optimization oracle with $\mathop{\mbox{poly}}(d,C,T)$ computational complexity in Section 4.
The key addition compared to (ito2020tight) is the bonus term $b_{t}$ that ensures misspecification robustness. This bonus encourages additional exploration, preventing the learner from being misled by misspecified feedback. The form of the bonus for adversarial linear bandits was first developed in the series of work (lee2020bias; zimmert2022return) aimed at obtaining high-probability bounds. Our use of the bonus is similar to liu2024towards, which tackles corruption and misspecification. In the regret analysis, this bonus term creates a negative regret term that offsets the regret overhead due to misspecification.
@@
We remark that while there exist algorithms that are $1$-misspecification-robust (liucorruption), their run time scales at least with the number of actions. Algorithm 2 achieves $\alpha$-robustness with the smallest $\alpha$ we are aware of among algorithms that run in $\mathop{\mbox{poly}}(d,C,T)$ time.
@@
1Parameters: $\gamma=10\log(10dT)$, $\beta=T^{-4}$.
@@ 3   Define
$q_{t}(y)=\frac{\exp\left(-\eta\sum_{\tau<t}\left\langle y,\widehat{\theta}_{\tau}-b_{\tau}\right\rangle\right)}{\int_{\widehat{\Omega}}\exp\left(-\eta\sum_{\tau<t}\left\langle z,\widehat{\theta}_{\tau}-b_{\tau}\right\rangle\right)\mathrm{d}z},\ \ x_{t}=\mathbb{E}_{y\sim q_{t}}[y],\ \ \Sigma_{t}=\mathbb{E}_{y\sim q_{t}}[(y-x_{t})(y-x_{t})^{\top}].$
@@ 4   Define
$\widehat{q}_{t}(y)=\frac{q_{t}(y)\mathbb{I}\left\{\|y-x_{t}\|_{\Sigma_{t}^{-1}}^{2}\leq d\gamma^{2}\right\}}{\int_{\widehat{\Omega}}q_{t}(z)\mathbb{I}\left\{\|z-x_{t}\|_{\Sigma_{t}^{-1}}^{2}\leq d\gamma^{2}\right\}\mathrm{d}z},\ \ \widehat{\Sigma}_{t}=\mathbb{E}_{y\sim\widehat{q}_{t}}[(y-x_{t})(y-x_{t})^{\top}].$
@@
6  [3pt] Define $\widehat{\theta}_{t}=(\beta I+\widehat{\Sigma}_{t})^{-1}(y_{t}-x_{t})c_{t}$ and $b_{t}=8\eta\left(\epsilon+\frac{1}{T^{2}}\right)\sum_{\tau<t}(\widehat{\theta}_{\tau}-b_{\tau})$.
@@ 3.5 Combining Everything and Using the Doubling Trick
Given access to a $\alpha$-misspecification-robust adversarial linear bandit algorithm $\mathsf{ALG}$, Algorithm 1 with doubling trick achieves $\max_{\pi\in\Pi}\mathrm{Reg}_{T}(\pi)\leq\widetilde{O}(d\sqrt{T}+\alpha d\sqrt{T\log K})$ in adversarial linear contextual bandits without access to simulators.
@@
$\max_{\pi\in\Pi}\mathrm{Reg}_{T}(\pi)\leq\tilde{O}\left(\sum_{k=1}^{\log_{2}T}\left(d\sqrt{2^{k}}+\alpha d\sqrt{2^{k}\log(NKT)}\right)\right)=\tilde{O}\left(d\sqrt{T}+\alpha d\sqrt{T\log K}\right).$
@@ 4 Computational Complexity
In this section we discuss how to construct the $\mathop{\mbox{poly}}(d,C,N)$ time linear optimization oracle needed in Section 3.4 to make Algorithm 2 efficient; we also consider two steps from Algorithm 1, lines 1 and 1, for which it is not obvious how they can be implemented efficiently. We will describe an approach to implement both steps in $\mathop{\mbox{poly}}(d,C,N)$ time, provided that $\operatorname{conv}(\mathcal{A})$ is a polytope that can be described by at most $C$ linear constraints for $\mathcal{D}$-almost all $\mathcal{A}$.
Constructing Linear Optimization and Separation Oracles.
By Corollary 14.1a of ilp_theory (restated in Lemma 4), we can obtain a $\mathop{\mbox{poly}}(d,C,N)$ time linear optimization oracle for $\widehat{\Omega}$ from a $\mathop{\mbox{poly}}(d,C,N)$ time separation oracle for $\widehat{\Omega}$. We proceed to construct such a separation oracle for $\widehat{\Omega}$. To this end, let us assume, without loss of generality, that no two sets $\mathcal{A}_{i}$ and $\mathcal{A}_{j}$ in the construction of $\widehat{\Omega}$ are equal; otherwise we can replace them by a single set $2\mathcal{A}_{i}$, which will only decrease $N$. Thus, $\widehat{\Omega}$ is the Minkowski sum of $N$ convex polytopes that, by assumption, can all be described by at most $C$ constraints. For any $x\in\mathbb{R}^{d}$, let
@@
If $g(\phi)>0$ for some $\phi$, then $\phi$ gives a hyperplane that separates $x$ from $\widehat{\Omega}$; and if $g(\phi)\leq 0$ for all $\phi$, then $x\in\widehat{\Omega}$. Since $g$ is concave, and we can solve every subproblem $\max_{x_{i}\in\mathcal{A}_{i}}\phi^{\top}x_{i}$ in $\mathop{\mbox{poly}}(d,C)$ time, we can maximize $g$ in $\mathop{\mbox{poly}}(d,C,N)$ time to obtain our separation oracle.
Implementing line 1 of Algorithm 1.
@@
To implement line 1, we need to be able to take any point $y_{t}\in\widehat{\Omega}$ and find a distribution $\alpha_{t}$ supported on the vertices of $\widehat{\Omega}$ that we can sample from efficiently. By Carathéodory’s theorem, any $y_{t}\in\widehat{\Omega}$ can be represented as a convex combination of at most $d+1$ vertices of $\widehat{\Omega}$: $y_{t}=\sum_{l=1}^{k}\alpha_{l}v_{l}$ for $k\leq d+1$ where all $v_{l}$ are vertices of $\widehat{\Omega}$, and $\alpha_{1},\dots,\alpha_{k}\geq 0$ with $\sum_{l}\alpha_{l}=1$ can be interpreted as the probabilities of selecting the vertices. This is a categorical distribution on at most $d+1$ points, from which we can sample in $O(d)$ time (assuming we have access to an oracle that provides samples from the uniform distribution on $[0,1]$). Thus the main challenge is to compute the vertices $v_{l}$ and the probabilities $\alpha_{l}$. By Corollary 14.1g) of ilp_theory (restated in Lemma 3), there exists an algorithm that can do both in time $\mathop{\mbox{poly}}(d,h)$ for any set $\widehat{\Omega}$ given access to a separation oracle that runs in time $O(h)$. As discussed above, we have $h$ of order $\mathop{\mbox{poly}}(d,C,N)$, from which it follows that we can also find the probabilities $\alpha_{l}$ and vertices $v_{l}$ in $\mathop{\mbox{poly}}(d,C,N)$ time.
Implementing line 1 of Algorithm 1.
After the sampling procedure has chosen a particular vertex $\psi_{t}\in\{v_{1},\ldots,v_{k}\}$, we need to find an interior point of the corresponding normal cone $\mathcal{N}(\widehat{\Omega},\psi_{t})$ in line 1. To achieve this, suppose we are given a vertex $\psi_{t}\in\widehat{\Omega}$. By Lemma 5, there exists a unique decomposition $\psi_{t}=\frac{1}{N}\sum_{i=1}^{N}x_{i}$ where $x_{i}$ is a vertex of $\operatorname{conv}(\mathcal{A}_{i})$ for all $i\in[N]$. We can find $\{x_{i}\}_{i\in[N]}$ by solving a linear program with $O(NC)$ constraints: $x_{i}\in\operatorname{conv}(\mathcal{A}_{i})$ for all $i\in[N]$ and $\frac{1}{N}\sum_{i=1}^{N}x_{i}=\psi_{t}$.
Let $\mathcal{N}_{i}=\mathcal{N}(\operatorname{conv}(\mathcal{A}_{i}),x_{i})$ be the normal cone of $\operatorname{conv}(\mathcal{A}_{i})$ at the vertex $x_{i}$. It is the conic hull of the active constraint normals: $\mathcal{N}_{i}=\operatorname{cone}(\mathcal{T}_{i})=\left\{\sum_{u\in\mathcal{T}_{i}}\lambda_{i,u}u~\big|~\lambda_{i,u}\geq 0\text{\ for all }u\in\mathcal{T}_{i}\right\}$, where $\mathcal{T}_{i}$ is the set of normal vectors $u$ corresponding to the constraints $u^{\top}x+v\leq 0$ that are active at $x_{i}$. $\mathcal{T}_{i}$ can be constructed by enumerating all constraints of $\operatorname{conv}(\mathcal{A}_{i})$ and checking which are tight at $x_{i}$.
By Lemma 6, $\mathcal{N}(\widehat{\Omega},\psi_{t})=\bigcap_{i\in[N]}\mathcal{N}_{i}$. To find a point strictly inside $\mathcal{N}(\widehat{\Omega},\psi_{t})$, we solve the following linear program with $O(NC)$ constraints:
$\max_{\varepsilon\in\mathbb{R},\ \lambda_{i,u}\in\mathbb{R},\ \phi\in\mathbb{R}^{d}}\ \varepsilon$
$\text{subject to}\ \ \phi=-\sum_{u\in\mathcal{T}_{i}}\lambda_{i,u}u\quad\text{for all }i\in[N],$ ($\phi\in-\mathcal{N}_{i}$ for all $i$)
$\qquad\qquad\ \varepsilon\leq\lambda_{i,u}\leq 1\quad\text{ for all }i\in[N]\text{ and }u\in\mathcal{T}_{i}.$ (Make $\phi$ bounded and strictly inside $\mathcal{N}_{i}$)
By the assumption that $\psi_{t}$ is a vertex, $\mathcal{N}(\widehat{\Omega},\psi_{t})$ always has an non-empty interior. Therefore, a solution with $\varepsilon>0$ exists, and the optimal $\phi$ provides a desired interior point.
@@ 5 Small-Loss Bound with Access to Simulator
A misspecification-robust linear bandit algorithm with small-loss bounds over action set $\widehat{\Omega}\subset\mathbb{R}^{d}$ has the following property: with a given $\epsilon>0$ and the guarantee that every time the learner chooses $y_{t}\in\widehat{\Omega}$, the loss received $c_{t}\in[0,1]$ satisfies $|\mathbb{E}_{t}[c_{t}]-\left\langle y_{t},\theta_{t}\right\rangle|\leq\epsilon$, the algorithm ensures
@@ 6 Conclusion and Open Questions
Three open questions are left by our work: First, can the computation cost be improved further to match that of neu2014online, who only require a linear optimization oracle for each action set, and do not require a polynomial number of constraints? Second, can the regret be further improved to the near-optimal $\widetilde{\mathcal{O}}(d\sqrt{T})$ bound with a polynomial-time algorithm without simulators? Third, can we achieve $\mathop{\mbox{poly}}(d)\sqrt{L^{\star}}$ regret without simulator? For the second question, one may try to generalize the approach of schneider2023optimal. For the third question, one idea is to establish a Bernstein-type counterpart of Lemma 1, potentially drawing ideas from bartlett2005local; liang2015learning.
We expect that our approach has wider applications than adversarial linear contextual bandits. For example, our approach may be generalized to linear MDPs with fixed transition and adversarial losses (luo2021policy; sherman2023improved; dai2023refined; kongimproved; liu2024towards) and facilitate learning with exponentially large or continuous action sets.
Acknowledgments and Disclosure of Funding
Mayo and Van Erven were supported by the Netherlands Organization for Scientific Research (NWO) under grant number VI.Vidi.192.095.
Appendix A Comparison with hanna2023contexts
Our work and hanna2023contexts both reduce linear contextual bandits to linear bandits with misspecification. The linear bandits the two works reduce to are slightly different. For comparison, let $\mathcal{A}_{1},\ldots,\mathcal{A}_{N}$ be empirical action sets drawn from $\mathcal{D}$, and define
@@ Appendix A Comparison with hanna2023contexts
It holds that $\widehat{\Omega}_{\text{lin}}^{\epsilon}\subset\widehat{\Omega}_{\text{lin}}\subset\widehat{\Omega}$ because $\Pi^{\epsilon}_{\text{lin}}\subset\Pi_{\text{lin}}\subset\Pi$. The relation among the three sets in (6) is the following: $\widehat{\Omega}_{\text{lin}}$ is the boundary of $\widehat{\Omega}$ (Lemma 2), and $\widehat{\Omega}_{\text{lin}}^{\epsilon}$ is a discretized subset of $\widehat{\Omega}_{\text{lin}}$. It has been shown that there must exist an optimal policy in $\Pi_{\text{lin}}$ (see our arguments in Section 2.1), and that $\Pi_{\text{lin}}^{\epsilon}$ must contain an $\epsilon$-near-optimal policy (see (47) of liu2024bypassing).
Our work reduces the linear contextual bandit problem to linear bandits in $\widehat{\Omega}$, while hanna2023contexts reduces it to linear bandits in $\widehat{\Omega}^{\epsilon}_{\text{lin}}$. According to the discussion above, both suffice to establish no regret guarantees against the optimal policy. The key difference lies in the computational efficiency. As discussed in Section 4, $\widehat{\Omega}$ is a polytope whose separation oracle can be implemented in $\mathop{\mbox{poly}}(d,C,T)$ time, while in hanna2023contexts, $\widehat{\Omega}^{\epsilon}_{\text{lin}}$ is a set containing $(\frac{1}{\epsilon})^{\Theta(d)}$ discrete points with $\epsilon$ chosen as $\Theta(\frac{1}{T})$. It remains open whether $\widehat{\Omega}^{\epsilon}_{\text{lin}}$ can be represented efficiently and whether hanna2023contexts’s algorithm can be implemented in polynomial time. In addition, our work operates on the adversarial-loss setting, strictly generalizing hanna2023contexts’s stochastic-loss setting.
A technical question related to the discussion above is whether we can apply the discretization idea from hanna2023contexts purely in the analysis but not in the algorithm, i.e., bounding the regret overhead by the discretization error without restricting the learner’s policy to the discretized linear policies. This would allow us to avoid the Rademacher complexity analysis (Lemma 1) and instead use Hoeffding’s or Bernstein’s inequality combined with a union bound, possibly leading to an improved regret bound. However, there is a key technical challenge: the mapping $\hat{\Psi}(\pi_{\phi})$ is not continuous in $\phi$, which prevents us from bounding $\|\hat{\Psi}(\pi_{\phi})-\hat{\Psi}(\pi_{\phi^{\prime}})\|$ by a constant times $\|\phi-\phi^{\prime}\|$, where $\phi$ is an arbitrary point and $\phi^{\prime}$ is a point in the discretized set. This would leave us no control on the performance of policies outside the set of discretized linear policies $\Pi^{\epsilon}_{\text{lin}}$. As a result, we are unable to derive a uniform convergence result analogous to Lemma 1 using this approach.